# Information Extraction From Fiscal Documents Using LLMs

### v1

@@ 5. Future Directions
We have demonstrated a technique for extracting information from state fiscal documents by using a single Indian state (Karnataka) as an example. Our research goal is to extend this work across all Indian states and over many years of fiscal disclosures, to create a robust, analyzable dataset of state finances. To achieve this goal, here are the future directions for our work:
@@
    • Generates CSV schemas per table type.
@@
5.1. Application Areas
While this paper has a narrow focus: the task of extracting information from state finances; Our overarching goal is to use this data to understand and improve the functioning of states. This work will find value in:
  • Public procurement analysis
  • Budgeting efficiency studies (Budgeted vs Revised vs Actuals)
  • Expenditure priority tracking over time
  • In-year fund reallocation analysis
  • Functional/departmental/geographic classification
@@ 5.2. Reproducible Research
All materials, including the fiscal documents, source code, validation scripts, extracted information and instructions are available under MIT license at: https://www.github.com/xkdr/acm_icaif_2025_paper.
@@ 6. Conclusion
We have demonstrated that LLMs can successfully read and understand complex hierarchical tables when provided with appropriate domain knowledge, sequential context, and validation mechanisms. Our approach understands the structure of 200+ page PDF documents, and extracts it into research-ready machine-readable files with a numerical accuracy of 84%, and structural accuracy ranging from 73% to 96%. Our mechanism also provides the locations of data extraction failures, creating a technique that can be used to generate verifiable seed datasets for downstream analysis.
The key insight is that LLMs bring semantic understanding to table extraction – they don’t just perform pattern matching but leverage domain knowledge acquired from large-scale pretraining. Combined with built-in validation through summation checks, this enables reliable automation of a task previously requiring extensive manual effort.
While it has been applied here to Indian state finances, the methodology is broadly applicable where complex tabular data is published in non-machine-readable formats and where internal consistency checks can be defined. This opens up new avenues for leveraging LLMs in public finance research, transparency initiatives, and data democratization efforts, but also in generic table extraction tasks across domains.

### v2

@@ 5. Future Directions
We have demonstrated a technique for extracting information from state fiscal documents by using a single Indian state (Karnataka) as an example. Our research goal is to extend this work across all Indian states and over many years of fiscal disclosures, to create a trusted database about Indian state finance. This will require many improvements in the implementation:
@@
    • Generates CSV schema per table type.
@@
  • Create large scale reference datasets of table databases that can feed back into AI training for the future.
@@ 5.1. Reproducible Research
All materials, including the fiscal documents, source code, validation scripts, extracted information and instructions are available under MIT license at: https://github.com/xKDR/LLM_table2db.
@@ 6. Conclusion
We have shown a path for LLMs to read and understand complex hierarchical tables when provided with appropriate domain knowledge, sequential context, and validation mechanisms. Our approach understands the structure of 200+ page PDF documents, and extracts it into research-ready machine-readable files with high accuracy. Our mechanism also provides the locations of data extraction failures, even though the ground truth is not observed.
These results suggest that LLMs do much more than regular expression matching. They are able to leverage their vast domain knowledge acquired in pre-training, so that prompts are able to steer them effectively. Combined with built-in validation through summation checks, this enables reliable automation of a task previously requiring extensive manual effort.
While it has been applied here to Indian state finances, the methodology is broadly applicable where complex tabular data is published in non-machine-readable formats and where internal consistency checks can be defined. This opens up new avenues for leveraging LLMs for extracting datasets from PDF files in many situations.